### v1

# Nicer Than Humans: How do Large Language Models Behave in the Prisoner’s Dilemma?

Abstract
The behavior of Large Language Models (LLMs) as artificial social agents is largely unexplored, and we still lack extensive evidence of how these agents react to simple social stimuli. Testing the behavior of AI agents in classic Game Theory experiments provides a promising theoretical framework for evaluating the norms and values of these agents in archetypal social situations. In this work, we investigate the cooperative behavior of Llama2 when playing the Iterated Prisoner’s Dilemma against random adversaries displaying various levels of hostility. We introduce a systematic methodology to evaluate an LLM’s comprehension of the game’s rules and its capability to parse historical gameplay logs for decision-making. We conducted simulations of games lasting for 100 rounds, and analyzed the LLM’s decisions in terms of dimensions defined in behavioral economics literature. We find that Llama2 tends not to initiate defection but it adopts a cautious approach towards cooperation, sharply shifting towards a behavior that is both forgiving and non-retaliatory only when the opponent reduces its rate of defection below 30%. In comparison to prior research on human participants, Llama2 exhibits a greater inclination towards cooperative behavior. Our systematic approach to the study of LLMs in game theoretical scenarios is a step towards using these simulations to inform practices of LLM auditing and alignment.
1 Introduction
Large Language Models (LLMs) can operate as social agents capable of complex, human-like interactions (Park et al. 2023). Their integration into online social platforms is unfolding rapidly (Cao et al. 2023; Yang and Menczer 2023), presenting severe risks (Floridi and Chiriatti 2020; Ferrara 2024) as well as intriguing opportunities (Dafoe et al. 2020; Breum et al. 2023). To understand and anticipate the behavioral dynamics that may arise from the interaction between artificial agents and humans, it is essential to first study how these agents react to simple social stimuli (Bail 2024). Behavioral economics experiments, particularly those grounded in Game Theory, provide an ideal ground for testing the responses of AI agents to archetypal social situations (Horton 2023). These experiments typically involve goal-oriented scenarios where multiple ‘players’ engage in a series of repeated interactions (Osborne and Rubinstein 1994). To optimize for the goal, the decisions taken at each round must strategically account for the anticipated actions of the other players. However, the decisions of human participants often deviate from the theoretically optimal strategies due to the influence of social and psychological factors that conflict with the game’s objectives (Camerer 1997). Similarly, given that LLMs are computational models built upon collective human knowledge and culture (Schramowski et al. 2022), observing their behavior in classic iterated games could shed light on the social norms and values that these models reflect, as well as their capability in reasoning, planning, and collaborating in social settings.
Early interdisciplinary research has explored the use of LLMs within the context of classical economic games (see Related Work). While highly valuable, all these studies exhibit at least one of the following limitations. First, they generally lack prompt validation procedures, leading to an implicit assumption that LLMs can understand the complex rules of the game and the history of past actions described in the prompt (Akata et al. 2023; Mao et al. 2023; Mei et al. 2024). Second, the duration of simulated games is often limited to a few rounds (Akata et al. 2023; Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Xu et al. 2023), hampering the LLMs’ ability to discern the decision-making patterns of other participants – a phenomenon we show in our own experiments. Third, the initialization of LLMs with predefined ‘personas’ tends to skew their responses towards pre-determined behaviors, such as altruism or selfishness (Brookins and DeBacker 2023; Fan et al. 2023; Guo 2023; Horton 2023; Lorè and Heydari 2023; Phelps and Russell 2023). This approach limits the exploration of the LLMs’ baseline behavior, which is crucial for understanding their inherent decision-making processes. Fourth, the evaluation of simulation outcomes has predominantly focused on the quantitative analysis of decision types (e.g., frequency of cooperation), overlooking the LLMs’ higher-level behavioral patterns that can be inferred from the temporal evolution of these decisions (Xu et al. 2023; Mao et al. 2023). The combined effect of these limitations has led to findings that are sometimes inconclusive (Brookins and DeBacker 2023; Mao et al. 2023) and contradictory (Akata et al. 2023; Fan et al. 2023), calling for more systematic evidence on the behavior of LLMs in iterated games.
In this work, we investigate the adaptability of LLMs in terms of their cooperative behavior when facing a spectrum of hostility in an iterated game scenario. We evaluate Llama2 (Touvron et al. 2023) in the Iterated Prisoner’s Dilemma (Osborne and Rubinstein 1994) against adversaries with different propensities for betrayal. Our contribution is threefold. First, we introduce a meta-prompting technique designed to evaluate an LLM’s comprehension of the game’s rules and its ability to parse historical gameplay logs for decision-making. Second, we conduct extensive simulations over 100 rounds and determine the optimal memory span that enables the LLM to adhere to the strategic framework of the game. Third, we analyze the behavioral patterns exhibited by the LLM, aligning them with the core dimensions and strategies delineated in Robert Axelrod’s influential research on the evolution of cooperation within strategic interactions (Axelrod and Hamilton 1981).
We find that Llama2’s behavior is generally ‘nice’, as it tends not to defect first. It adopts a conservative stance towards cooperation, sharply shifting towards behavior that is both forgiving and less retaliatory when the opponent reduces its rate of defection under 30%. Overall, Llama2 displayed a more marked propensity towards cooperation than what existing literature reports about human players.
Overall, our work contributes to defining a more principled approach to using LLMs for iterated games. It makes a step towards a more systematic way to use simulations of game theoretical scenarios to probe the inherent social biases of LLMs, which might prove useful for LLM auditing and alignment (Shen et al. 2023; Mökander et al. 2023).
2 Background on Prisoner’s Dilemma
2.1 Game Setup
The Prisoner’s Dilemma is a classic thought experiment in Game Theory. It serves as a paradigm for analyzing conflict and cooperation between two players (Tucker and Straffin Jr 1983). In the game, the two players cannot communicate, and must independently choose between two actions: Cooperate, or Defect. Once both players have chosen their actions, payoffs are distributed based on the resulting combination of their choices. Mutual cooperation yields a reward $R$ for each player. If one defects while the other cooperates, the defector receives a higher ‘temptation’ payoff $T$, while the cooperating player incurs a lower ‘sucker’s’ payoff $S$. If both parties choose to defect, they each receive a punishment payoff $P$ for failing to cooperate. The classical structure of the game is defined by the payoff hierarchy $T>R>P>S$, which theoretically incentivizes rational players to consistently choose defection as their dominant strategy (Axelrod 1981). In the iterated version of the game, multiple rounds are played, and the payoffs are revealed to the players at every round (Tucker and Luce 1959). The iterative nature of the game allows the players to consider past outcomes to strategically inform future actions. When humans play the game, psychological factors such as reputation and trust significantly influence the decision-making process, often leading to higher rates of cooperation than would be expected from purely rational agents (Dal Bó and Fréchette 2011; Romero and Rosokha 2018).
2.2 Strategies
In the Iterated Prisoner’s Dilemma (IPD), a strategy refers to an algorithm that a player uses to decide their next action, taking into account the historical context of the game (Kuhn 2019). No single strategy universally outperforms all others; however, some are more effective against a broader variety of opposing strategies (Dal Bó and Fréchette 2011). This concept was demonstrated in 1980 by Robert Axelrod (Axelrod 1980), who ran an IPD tournament with multiple competing strategies. Follow-up tournaments have further diversified the range of strategies (Stewart and Plotkin 2012). Considering previous literature (Fudenberg, Rand, and Dreber 2012; Dal Bó and Fréchette 2011; Romero and Rosokha 2018), we consider the strategies that better represent the ones adopted by humans, covering more than $75\%$ of the experimental samples in those studies. Those strategies are the following:
  • $\mathsf{Always\ Cooperate}$ ($\mathsf{AC}$).
  • $\mathsf{Always\ Defect}$ ($\mathsf{AD}$).
  • $\mathsf{Random}$ ($\mathsf{RND}$). Chooses Cooperate or Defect at random with equal probability at each round.
  • $\mathsf{Unfair\ Random}$ ($\mathsf{URND}p$). Variation of Random where the probability of choosing to Cooperate is $p$.
  • $\mathsf{Tit\ For\ Tat}$ ($\mathsf{TFT}$). Starts with Cooperation in the first round, then mimics the opponent’s previous action throughout the game.
  • $\mathsf{Suspicious\ Tit\ For\ Tat}$ ($\mathsf{STFT}$). A $\mathsf{TFT}$ strategy that begins with Defect in the first round.
  • $\mathsf{Grim\ Trigger}$ ($\mathsf{GRIM}$). Chooses Cooperate until the opponent defects, then chooses only Defect for the rest of the game.
  • $\mathsf{Win\mbox{--}Stay\ Lose\mbox{--}Shift}$ ($\mathsf{WSLS}$). Repeats the previous action if it resulted in the highest payoffs ($R$ or $T$), otherwise changes action.
$\mathsf{Tit\ For\ Tat}$ emerged as the winning strategy in Axelrod’s tournament. It is commonly observed that human players tend to favor straightforward strategies such as $\mathsf{AD}$, $\mathsf{TFT}$, or $\mathsf{GRIM}$ (Romero and Rosokha 2018). To describe the LLM behavior in terms of these known strategies, in our experiments we calculate the similarity of the LLM’s game sequences with the sequences that these hardcoded strategies would generate when playing against the same opponent.
2.3 Behavioral Dimensions
To identify the defining factors of different strategies, we combine dimensions already defined in prior studies that quantify salient behavioral properties of players based on the observed game sequences (Axelrod 1980; Mei et al. 2024):
  • $\mathsf{Nice}$. Propensity to not be the first to defect. For a single instance of the Iterated Prisoner’s Dilemma, it is defined as $1$ if the player is not the first to defect, 00 otherwise.
  • $\mathsf{Forgiving}$. Propensity to cooperate again after an opponent’s defection, defined as: #$\frac{\#\mathsf{forgiven\_defection}}{\#\mathsf{opponent\_defection}+\#\mathsf{penalties}},$ where the number of penalties corresponds to the times that, after defecting, the opponent sought forgiveness by cooperating and the player did not forgive them, thus keeping defecting.
  • $\mathsf{Retaliatory}$. Propensity of defecting immediately after an opponent’s uncalled defection, defined as: #$\frac{\#\mathsf{reactions}}{\#\mathsf{provocations}}$
  • $\mathsf{Troublemaking}$. Propensity to defect unprovoked, defined as a counterpart of being retaliatory: #$\frac{\#\mathsf{uncalled\_defection}}{\#\mathsf{occasions\_to\_provoke}}$, where an uncalled defection is a defection following a cooperation (or being the first action of the game) and an occasion to provoke is any cooperation from the opponent in the previous round.
  • $\mathsf{Emulative}$. Propensity to copy the opponent’s last move: #$\frac{\#\mathsf{mimic}}{N-1}$, where a mimic occurs any time the player played the same action that the opponent played in the previous round and $N$ is the number of iterations of the game.
Strategies that are $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ (e.g., $\mathsf{TFT}$) perform best against a wide variety of opponents. Human players tend to be particularly uncooperative when faced with games where the reward $R$ for cooperating is much lower than the temptation $T$ to betray the other player. In the indefinitely iterated version of the Prisoner’s Dilemma with a fixed probability at every round for the game to terminate, human subjects tend to be more cooperative when the probability of ending the game is low. Overall, in games with a low chance of continuation and a big gap between $R$ and $T$, the majority of the strategies adopted are most similar to $\mathsf{Always\ Defect}$ (70% to 90%); in games with a longer potential time horizon and $R$ closer to $T$, humans tend to be more forgiving and $\mathsf{Tit\ For\ Tat}$ explains a larger portion of the subjects’ strategies (Dal Bó and Fréchette 2011).
3 Experimental Design
3.1 LLM Setup
For our experiments, we use the Llama-2-70B-chat model, an open-source LLM, released under a commercial use license [^1] https://ai.meta.com/llama/license/. We use a version of the model hosted by Hugging Face, which we access through the Inference API [^2] https://huggingface.co/inference-api/serverless. We initialize Llama2 with a temperature of $0.7$, in line with the setup of previous work (Lorè and Heydari 2023; Xu et al. 2023).
We make Llama2 play a series of Iterated Prisoner’s Dilemma games, each consisting of $N=100$ rounds. Due to the stochastic nature of the responses that the LLM generates, we repeat each game $k=100$ times and report the average results along with 95% confidence intervals. To evaluate the LLM’s adaptability to different degrees of environmental hostility, we repeat the experiment by matching the LLM against $\mathsf{URND}p$ opponents (defined in §2.2) with varying probability of cooperation $\alpha\in[0,1]$. The final outcome of each game is a sequence containing pairs of binary values representing representing cooperation (1) or defection (0) by the LLM (player $A$) and the opponent (player $B$) at each round $i$:
$G^{\alpha}_{k}=[(A_{i},B_{i})]_{i\in[1,N]}.$ (1)
From the $G^{\alpha}_{k}$ sequence, we extract the empirical probability of the LLM to cooperate at round $i$, calculated as the fraction of $i^{th}$ rounds in which the LLM cooperated over $k$ trials:
$p^{\alpha}_{coop}(i)=\frac{1}{k}\sum_{k}{G^{\alpha}_{k}(A_{i})}.$ (2)
We calculate the average cooperation probability throughout a game by averaging $p^{\alpha}_{coop}(i)$ over all $N$ rounds:
$p^{\alpha}_{coop}=\frac{1}{N}\sum_{i=1}^{N}p^{\alpha}_{coop}(i).$ (3)
3.2 Prompting
To implement the game, we design a fixed system prompt that outlines the game’s rules, including the payoff structure, and the player’s objective to ‘get the highest possible number of points in the long run’. The variable part of the prompt includes the memory of the game, namely a log of the history of the players’ actions up to the current round, along with instructions for generating the next action. The complete prompt can be found in the Appendix. In formulating the memory component, we explore various window sizes to provide the LLM with information from only the $n$ most recent rounds. We determine the optimal window size through targeted experiments that matched the LLM against $\mathsf{AC}$, $\mathsf{AD}$, and $\mathsf{RND}$ players (see §4.2).
3.3 Meta-prompting
The development of effective LLM prompts is an ever-evolving practice. Although certain studies suggested guidelines for prompt development (Ziems et al. 2023), achieving a consensus on the most effective prompting strategies across tasks remains challenging. Typically, the prompt quality is assessed empirically based on downstream performance (Lester, Al-Rfou, and Constant 2021). This method is suitable for conventional classification or regression tasks where some form of ground truth is clearly defined. However, it is less applicable to generative tasks that lack a formal standard of correctness. In the specific context of the Prisoner’s Dilemma, any sequence of Cooperate and Defect actions could be considered plausible. This ambiguity makes it difficult to discern whether LLM-generated sequences reflect a proper understanding of the game’s rules or are merely the product of the model hallucinating (Xu, Jain, and Kankanhalli 2024). Prior research involving LLMs in Game Theory experiments has assessed outputs by requesting that the LLM provide a reasoned explanation of its output (Guo 2023). However, this approach offers only a retrospective justification, which can itself suffer from hallucinations if the LLM has not fully grasped the task’s underlying instructions.
Name Question
Rules min_max What is the lowest/highest payoff player A can get in a single round?
actions Which actions is player A allowed to play?
payoff Which is player X’s payoff in a single round if $X$ plays $p$ and $Y$ plays $q$?
Time round Which is the current round of the game?
actioni Which action did player $X$ play in round $i$?
pointsi How many points did player $X$ collect in round $i$?
State #actions How many times did player $X$ choose $p$?
#points What is player $X$’s current total payoff?
Table 1: Templates of prompt comprehension questions used in meta-prompting to verify the LLM’s comprehension of the prompt.
To partially mitigate this issue, we introduce a novel meta-prompting technique to measure the LLM’s comprehension of a given prompt, to inform the process of prompt refinement. Specifically, we formulate a set of prompt comprehension questions that address three key aspects of the prompt (see Table 1): the game rules, to verify the LLM’s grasp of the game mechanics (e.g., ‘What is the lowest payoff that player A can get in a single round?’); the chronological sequence of actions within the game history (e.g., ‘Which action did player A play in round 5?’); and the cumulative game statistics (e.g., ‘What is player’s $B$ current total payoff?’).
To assess the LLM’s proficiency in responding to meta-prompting questions, we conduct three 100-round games against $\mathsf{RND}$ opponents. We pose the questions at each round and compute the average accuracy of the LLM’s responses. At any given round $i$, a question template is instantiated into a set of questions that cover all possible combinations of the template’s parameters. For example, at round $i$, questions referring to specific rounds (actioni and pointsi) are asked for all past rounds from 1 to $i-1$.
3.4 Behavioral Profiling
We profile the LLM’s behavior with respect to a game history $G^{\alpha}_{k}$ in two ways. First, we quantify behavioral patterns by computing the behavioral dimensions outlined in §2.3. This computation results in a five-dimensional numerical vector that encapsulates the behavioral characteristics of the LLMs. Second, we use the Strategy Frequency Estimation Method (SFEM) defined in previous work (Romero and Rosokha 2018) to calculate the affinity between a player’s game history and any of the classic strategies used in Prisoner Dilemma tournaments (see §2.2). SFEM is a finite-mixture approach that uses likelihood maximization to estimate the likelihood of a strategy being represented in experimental data (Romero and Rosokha 2018). Given a game history, SFEM outputs a score between 00 and $1$ for each strategy in the set of theoretical strategies considered. Being a mixture approach, the sum of all the SFEM scores over the set of possible strategies does not have to sum up to $1$. The final SFEM scores we report are averaged over all the histories analyzed.
4 Results
4.1 LLM’s Prompt Comprehension
Figure 1: Accuracy of Llama2 responses to the prompt comprehension questions defined in Table 1. Questions belong to three groups testing respectively aspects related to the rules of the game, its temporal evolution, and its current state. The accuracy of the final full prompt is compared with an earlier version of the prompt that lacked a summary of the cumulative rewards achieved by the players. 95% confidence intervals calculated over 100 rounds are shown.
Figure 1 presents the accuracy of responses to the prompt comprehension questions associated with the final prompt that we used in our experiments, with results averaged over all trial runs. Overall, Llama2 exhibits a good understanding of the concepts assessed by the questions, with most responses achieving an accuracy ranging from $0.8$ to $1.0$.
The process leading to finalizing the prompts was highly iterative, evaluating the accuracy of questions at each incremental step. Generally, adding explicit information about the game state and rules led to a better level or prompt comprehension. For illustration, Figure 1 includes a comparative analysis of accuracy scores derived from an earlier version of the prompt, which lacked a summary of the cumulative point totals for the players. In absence of explicit total counts, Llama2 was required to sum all points from the game history to determine the total tally, which significantly impacted its performance on the #actions and #points questions. This limitation of Llama2 aligns with previous findings that highlight that LLMs tend to struggle with arithmetic (Xu et al. 2023; Aher, Arriaga, and Kalai 2023; Wei et al. 2023). The explicit inclusion of the sum of scores into the prompt markedly improved performance, achieving near-perfect accuracy.
4.2 Effect of Memory Window Size
Figure 2: Left: Llama2’s probability of cooperation ($p_{coop}$) against an $\mathsf{Always\ Defect}$ opponent, when using a memory window size of 10 vs. including the full game history in the prompt. Right: steady-state probability calculated on the last 10 rounds of $p_{coop}$ for different memory windows sizes. 95% confidence intervals calculated over 100 games are shown.
In iterated games, the information from earlier rounds is essential for a player to deduce the opponent’s strategy, and adapt accordingly. Early research involving LLMs in iterated games experimented with a limited number of rounds, which precluded any analysis of how the size of the memory window influences the agent’s behavior (Akata et al. 2023; Guo 2023; Xu et al. 2023). Here, we analyze the behavior of Llama2 when it is provided with a history of the $n$ most recent rounds. We assess the impact of varying memory window sizes by matching the LLM against an $\mathsf{Always\ Defect}$ opponent. This test relies on the assumption that once the LLM accumulates enough information to infer the consistently defecting behavior of the opponent, the LLM actions should converge towards defection.
Figure 2 (left) shows the probability of the LLM cooperating at each round in games of $100$ rounds, considering two memory window sizes: $m=100$ and $m=10$. Under both conditions, the LLM shifts towards a stance of consistent defection after approximately $5$ to $10$ rounds. Notably, this duration equals or exceeds the maximum number of rounds considered in previous studies. Without any constraints on the history length, the LLM’s cooperation level quickly starts rising back, converging to full cooperation after the $50^{th}$ round. This pattern may be attributed to a combination of two factors: Llama2’s intrinsic preference for positive constructs (Lorè and Heydari 2023) (favoring cooperation over defection) and its limited effectiveness in extracting actionable insights from long prompts (Xi et al. 2023). With a memory window restricted to the $10$ most recent rounds, the LLM retains a full-defection stance throughout the game. We replicated this experiment across a range of memory window sizes and determined their respective equilibrium states by calculating the average cooperation probability in the final $10$ rounds (from the $90^{th}$ to the $100^{th}$). As illustrated in Figure 2 (right), memory windows sizes around $10$ yield the expected outcome. Therefore, we select a window size of $m=10$ for the remainder of our experiments.
4.3 Behavioral Patterns
Figure 3: Llama2’s probability of cooperation ($p_{coop}$) against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. 95% confidence intervals calculated over 100 games are shown.
We examine the overall propensity for cooperation exhibited by the LLM across various degrees of environmental hostility. Figure 3 shows the relationship between the probability of cooperative behavior $p^{\alpha}_{coop}$ and the varying cooperation levels $\alpha$ of an $\mathsf{Unfair\ Random}$ opponent. The LLM’s inclination to engage in cooperative behavior does not exhibit a linear relationship with $\alpha$. Instead, it follows a sigmoidal curve, indicative of a rapid transition from a predominantly defecting strategy to a more cooperative attitude. The sigmoid is characterized by a relatively flat left tail, maintaining a stable probability of cooperation near $p_{coop}=0$ for $\alpha$ values up to $0.4$. The curve’s inflection point lies between $0.6$ and $0.7$ — well beyond $0.5$ — suggesting that the LLM adopts a cautious approach when interpreting the actions of its opponent, rather than simply mirroring them.
The probability of cooperation provides a macroscopic perspective on a player’s actions. However, to capture more nuanced strategic patterns that emerge during the match, we employed SFEM analysis (defined in §3.4) to estimate the similarity between the behavioral patterns exhibited by the LLM and those commonly observed in games involving human players. Figure 4 presents the SFEM profile for the LLM against increasing values of $\alpha$. When the probability of the adversary’s cooperation surpasses the $0.6$ threshold, there is a discernible shift in the LLM’s strategy from a $\mathsf{Grim\ Trigger}$ strategy to an $\mathsf{Always\ Cooperate}$ approach. The SFEM analysis further indicates that these strategies are the most representative of the LLM’s strategic behavior, with no other strategies being significantly indicative.
At a finer level of analysis, we characterize the behavior of the LLM along the dimensions outlined in Section 2.3 (Figure 5). When the parameter $\alpha$ is set to low values, the LLM exhibits highly uncooperative traits: it frequently retaliates following instances of betrayal, seldom reverts to cooperative behavior after defecting, and is often the initiator of unprovoked defections. As $\alpha$ increases to 0.5 and beyond, we observe a linear transition in these behavioral dimensions, trending towards a more cooperative pattern. The LLM is consistently $\mathsf{Nice}$ across conditions, almost never defecting first. In the scenario where the opponent never defects ($\alpha=1$), the LLM never defects throughout the whole game in $50\%$ of the runs. Interestingly, there is no scenario in which the LLM is simultaneously $\mathsf{Nice}$, $\mathsf{Forgiving}$, and $\mathsf{Retaliatory}$ — the three conditions that characterize the most successful strategies, such as $\mathsf{Tit\ For\ Tat}$.
Figure 4: SFEM scores quantifying the similarity between Llama2’s sequences of actions and known strategies adopted in the Iterated Prisoner’s Dilemma game (defined in §2.2). Llama2’s behavioral sequences come from games against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$.
Figure 5: Presence of behavioral traits in Llama2’s actions when playing against $\mathsf{Unfair\ Random}$ opponents with increasing cooperation probability $\alpha$. The values of the same traits calculated for a $\mathsf{Random}$ agent playing against $\mathsf{Unfair\ Random}$ opponents are reported. Some traits are not defined for extreme values of $\alpha$. 95% confidence intervals calculated over 100 games are shown.
5 Discussion and Conclusion
Our study contributes to the broad literature on behavioral studies of Large Language Models as artificial social agents. Specifically, we study the responses of Llama2 to the prototypical social scenario of the Prisoner’s Dilemma. Building upon prior studies that explored the application of LLMs in classic Game Theory experiments, our work introduces a more systematic experimental setup that incorporates quantitative checks to better align the LLM responses to the complex task description. We have shown that aspects like prompt comprehension, memory representation, and duration of the simulation play crucial roles, with the potential to significantly distort the experimental outcomes, if left unchecked. Our framework provides a quantitative guide for selecting the simulation variables and improving the prompt.
Our findings add a new benchmark to the body of work that explored the outcomes of iterated games, both among humans and among AI agents. In contrast to the behavioral patterns observed in humans playing the Prisoner’s Dilemma game (Dal Bó and Fréchette 2011), Llama2 displayed a more marked propensity towards cooperation. Under conditions that disincentivize cooperation, most human players adopt a stance of complete defection, whereas the LLM’s strategy, albeit mostly uncooperative, is characterized by an initial trust in the opponent’s cooperation ($\mathsf{Nice}$), reminiscent of the strategy known in Game Theory as $\mathsf{Grim\ Trigger}$. When the environment is more favorable to cooperative play, human strategies often resemble either $\mathsf{Grim\ Trigger}$ or $\mathsf{Tit\ For\ Tat}$, while Llama2 tends towards a consistently cooperative approach. Notably, Llama2’s shift from $\mathsf{Grim\ Trigger}$ to $\mathsf{Always\ Cooperate}$ occurs quite abruptly when the probability of the opponent’s defection drops below 30%.
Our results are in line with early experiments involving LLMs, which indicated a tendency for these models to cooperate in repeated games (Brookins and DeBacker 2023; Mei et al. 2024). However, the broader research in this area has yielded mixed outcomes (Akata et al. 2023; Phelps and Russell 2023). Distinct from previous studies, our findings are derived from extensive game simulations conducted over numerous rounds and benefit from an experimental framework that leverages quantitative checks for accuracy.
However, it is important to acknowledge the limitations of our work, which open the way for further research. First, our analysis was conducted using a single model, specifically Llama2, which, at the time of writing, is one of the top-performing open-source models available (Touvron et al. 2023). Nevertheless, the field is progressing at an unprecedented pace, with new models being introduced regularly. To determine whether the behavioral patterns observed in our study are consistent across different models, it is important to conduct comparative analyses with models that vary significantly in terms of parameter size and the volume of their training data. Second, our study’s scope was limited to assessing the LLM’s responses to random strategies, and with a fixed payoff structure. Exploring the LLM’s interactions with more sophisticated opponents would enable us to better delineate the boundaries of LLMs’ inferential abilities in social contexts, and to draw more detailed behavioral profiles under a broader spectrum of conditions that more comprehensively represent prototypical social scenarios. Last, the experimental framework of our study considers only a single LLM agent. Creating social groups of AI agents that interact through iterated games like the Prisoner’s Dilemma would open up a wealth of opportunities to study emergent behaviors in synthetic societies, an avenue of research that is increasingly recognized as fundamental for a proper understanding of how LLMs can affect human societies (Bail 2024). Last, despite the numerical guidelines we implemented to evaluate the quality of the prompt, our refinements of the prompt were not guided by any principle other than experience and instinct. Our attempt to use Chain-of-Thought (Kojima et al. 2022; Zhou et al. 2023) as a structured way to approach prompt revision resulted in prompts that did not improve performance in our prompt comprehension question inventory. In this respect, our work provides yet another example of how prompt engineering would benefit from supporting tools to constrain its highly discretionary nature.
Despite its limitations, our work has two main implications. From the theoretical perspective, it expands our knowledge of the inherent biases of LLMs in social situations, which is crucial to inform their deployment across different contexts. From the practical perspective, it provides a principled way to approach game theoretical simulations with LLMs. This constitutes a step towards using these simulations as reliable and reproducible tools that could be used as tests to verify LLM alignment to desired principles of social cooperation (Shen et al. 2023).
6 Related work
Next, we briefly review previous work using LLMs for social reasoning, the generation of human-like synthetic data, and simulations of human behavior.
6.1 LLMs as Agents
Argyle et al. (2023) instructed LLMs to answer surveys as if they belonged to specified socio-demographic groups. They showed a high similarity between the responses generated by the LLM and those provided by the demographic groups it was asked to emulate. LLMs impersonating human agents with different profiles were used to create synthetic social networks, with the objective of observing emergent social behavior, most notably opinion dynamics and information spreading (Chuang et al. 2023; De Marzo, Pietronero, and Garcia 2023; He, Wallis, and Rathje 2023). Park et al. (2023) developed a society with synthetic agents interacting with elements of a synthetic world, showed that those agents were able to adopt behaviors that are typical of humans without being directly prompted to do so. Using the same framework, Ren et al. (2024) showed that those agents were also able to build and spread social norms.
6.2 LLMs in Game Theory
Early work on the application of LLMs to Game Theory experiments touched upon both 1-time and iterated games. Single-iteration experiments offer limited insight into LLM behavior. Brookins and DeBacker (2023) showed that LLMs are more biased towards fairness and cooperation when compared to a human baseline sample. In contrast, Aher, Arriaga, and Kalai (2023) found that there is an overall alignment between the LLM-based agent behavior and the human ones. When focusing on iterated games, the spectrum of patterns that can be identified expands, allowing more refined analysis. For example, Akata et al. (2023) managed to identify that LLMs can be particularly unforgiving. Mei et al. (2024) discovered instead that the same models show a higher cooperation rate than compared to humans. Fan et al. (2023) exploited the iteration of games to check the level of the opponent’s strategy that the LLM was able to infer from the history of actions. They showed that the inference capability of the LLM is limited, calling for more systematic approaches to structure memory and prompts.
7 Ethical Considerations
The deployment of Large Language Models as AI social agents raises numerous ethical considerations that are currently the subject of intense scrutiny by the interdisciplinary research community. The extraordinary capabilities of these models to generate text have led several scientists to envision alarming scenarios in which the seamless integration of AI agents into the online social discourse may facilitate the dissemination of harmful content, the spread of misinformation, and the propagation of ‘semantic garbage’, ultimately damaging our societies (Floridi and Chiriatti 2020; Weidinger et al. 2022; Hendrycks, Mazeika, and Woodside 2023). As a result, any research exploring the characteristics of LLMs as social agents could, directly or indirectly, contribute knowledge that might be exploited to implement and deploy LLM-based technologies for malicious purposes. While recognizing this risk, we also believe that conducting research on LLM-based agents is essential to assess potential risks and to guide efforts aimed at developing strategies to mitigate them. Our study contributes positively to deepen our understanding of how LLMs react to social stimuli.
Even when deploying LLM-based agents for ethical purposes, trade-offs between the obtained benefit and the high level of power consumption required to run them should be carefully considered (Bender et al. 2021).
Acknowledgments
NF acknowledges the support from the Danish Data Science Academy through the DDSA Visit Grant (Grant ID: 2023-1856) and from Politecnico di Milano through the scholarship “Tesi all’estero a.a. 2023/24-Primo bando”. FP acknowledges support from the European Union (NextGenerationEU project PNRR-PE-AI FAIR) and the Italian Ministry of Education (PRIN PNRR grant CODE prot. P2022AKRZ9 and PRIN grant DEMON prot. 2022BAXSPY). LMA acknowledges the support from the Carlsberg Foundation through the COCOONS project (CF21-0432).
Appendix
Figures 6, 7 and 8 show examples of the three parts of the final prompt we used in our experiments.
System prompt <s>[INST] <<SYS>>
Context: Player A and player B are playing a multi-round game. At each turn player A and player B simultaneously perform one of the following actions: {"Cooperate", "Defect"}
The payoffs for each combination of chosen actions are the following: If A plays "Defect" and B plays "Defect", A collects 1 points and B collects 1 points. If A plays "Defect" and B plays "Cooperate", A collects 5 points and B collects 0 points. If A plays "Cooperate" and B plays "Defect", A collects 0 points and B collects 5 points. If A plays "Cooperate" and B plays "Cooperate", A collects 3 points and B collects 3 points.
They will play a total of 100 rounds of this game.
Remember that a player’s objective is to get the highest possible number of points in the long run.
<<SYS>>
Figure 6: System prompt describing the players’ roles, the rules of the games, the number of rounds to be played, and the player’s objective. This part of the prompt is fixed in all simulations.
Contextual prompt
The history of the game in the last 5 rounds is the following:
Round 2: A played "Cooperate" and B played "Defect" A collected 0 points and B collected 5 points. Round 3: A played "Defect" and B played "Defect" A collected 1 points and B collected 1 points. Round 4: A played "Cooperate" and B played "Defect" A collected 0 points and B collected 5 points. Round 5: A played "Defect" and B played "Cooperate" A collected 5 points and B collected 0 points. Round 6: A played "Defect" and B played "Defect" A collected 1 points and B collected 1 points.
In total, A chose "Cooperate" 2 times and chose "Defect" 3 times, B chose "Cooperate" 1 times and chose "Defect" 4 times. In total, A collected 7 points and B collected 12 points.
Current round: 7.
Figure 7: Contextual prompt containing information about: the game history in the last $n$ rounds ($5$ in this example), the overall amount of times each player chose each action, the overall amount of points collected by each player, and the current round. For each turn, the prompt contains: the action played by each player and the points collected by each player. This prompt changes at each round.
Instructing prompt
Remember to use only the following JSON format:
{"action": <ACTION_of_A>, "reason": <YOUR_REASON>}
Answer saying which action player A should play.
Remember to answer using the right format.[/INST]
Figure 8: Instructing prompt. The LLM is instructed on the nature and format of the answer. This part of the prompt component is replaced with prompt comprehension questions in the phase of prompt tuning.